The following are recent year summaries of balance sheet and income statement data (in millions) for Apple Inc. (appl) and Verizon Communications (VZ). Determine the missing amounts. Enter decreases as negative amounts.

ugem-lake&inprogr eBook Calculator Accounting Equation The following are recent year summaries of balance sheet and income statement data (in millions) fo Determine the missing amounts. Enter decreases as negative amounts. Apple Verizon Year 1: Assets $231,839 Liabilities Stockholders equity Increase (Decrease) in assets, liabilities, and stockholders equity during Year 2: $38,836 $58,640 50,832 7,808 Assets Liabilities $(14,852) Stockholders equity Year 2: Assets $232,708 Liabilities $171,124 12,298 Stockholders equity Check My Work All work saved.
Answer #1

Calculate following

Apple Verizon
Year 1
Assets 231839 235262+38836 = 274098
Liabilities 171124-50832 = 120292 220410+14852 = 235262
Stockholder's equity 231839-120292 = 111547 38836
Increase (decrease) assets or liabilities
and stockholder's equity during year 2
Assets 58640 274098-232708 = -41390
Liabilities 50832 -14852
Stockholder's equity 7808 38836-12298 = -26538
Year 2:
Assets 231839+58640 = 290479 232708
Liabilities 171124 232708-12298 = 220410
Stockholder's equity 290479-171124 = 119355 12298